OpenStudy (anonymous):

A subway train makes six stops of equal length during its 21 km run. The train is actually moving for 20 min of the trip. At most how long can the train remain at each station if the average speed for the trip, including stops, is to be at least 36 km/h.

OpenStudy (anonymous):

Interesting question! Let me work it out on paper...

OpenStudy (anonymous):

Average Speed = Distance / Time Let total time stopped be x hrs

OpenStudy (anonymous):

Let t be the total amount of time. We know that t = 20 min + x min Let v be the average velocity along the whole trip, we know that v =36km/h Let d be the total amount of distance traveled, we know that d = 21 km We also know that d = v*t Can you solve this? It's a simple algebraic equation but you have to watch out for unit conversions.

OpenStudy (anonymous):

i got 21=36(20+x) but i don't think this is correct

OpenStudy (anonymous):

You have written distance = speed * time and this is the same as speed = distance/time so now just solve for x...

OpenStudy (anonymous):

Right. But don't forget the units in your equations. It is very important!

OpenStudy (anonymous):

I solved it before but I threw my paper away so I can't check the solution for you, but I remember the unit conversions where the most tricky part, because it's easy to forget :)

OpenStudy (anonymous):

but when i solved it i got 21= 720 + 36x which is 699=36x so x would be 19.4166667

OpenStudy (anonymous):

Your time has to be in hours....

OpenStudy (anonymous):

Remember I said "Let total time stopped be x hrs"

OpenStudy (anonymous):

Or you couold convert your speed to km per minute but that's more complicated...

OpenStudy (anonymous):

20 minutes = 1/3 hours

OpenStudy (anonymous):

but i still don't understand how you would convert the whole thing. what would be the equation i would solve?

OpenStudy (anonymous):

you had the right equation but you put (20 + x) for the time. You have to put (1/3 + x)

OpenStudy (anonymous):

9= 36x x= 1/4 (15 min) but i looked in the book and the answer said at most 2 and 1/2 minutes.

OpenStudy (anonymous):

but i don't get how to get that answer

OpenStudy (anonymous):

x is the total time stopped ie all 6 stops and 15/6 = 5/3 or 1 and 2/3 minutes per stop I don't know how they are calculating the "average" I also don't know why they are saying "at most"

OpenStudy (anonymous):

but i still don't understand how you get the answer

OpenStudy (anonymous):

You had 21=36(20+x) but time should be in hours so 21 = 36(1/3 +x) = 12 + 36 x so x = 9/36 = 174 hour or 15 minutes...